# Characterization of coherent errors in gate layers with robustness to Pauli noise

### v2

@@ III Characterization
We conduct the characterization of the same device and gate layer as in Fig. 3 three times, with time differences between the runs being 6 hours and 19 days. The results available in App. E show that the leading coherent errors in the system do not change over time. The most significant single-qubit coherent errors change by no more than $10\%$ and the two-qubit coherent rotation errors by maximally $0.018$ radians. This result suggests that the platform suffers from systematic coherent errors not influenced by the daily re-calibration of the system.
@@ IV Mitigation
In principle, the characterization results in Fig. 3 may originate from overfitting the non-ideal process to the introduced model. Indeed, it is unknown whether the coherently rotated Pauli noise model in Eq. (2) is expressive enough to resemble the main characteristics of the probed IBMq process. To verify that the model captures the dominant noise contributions of a process on the probed device, we examine whether the output of the presented protocol facilitates significant noise mitigation
@@
Returning to the unmitigated circuits in Fig. 1, our goal is to suppress the observed oscillatory behavior in the echo-experiment by correcting the characterized coherent errors at the circuit level. In case of a successful correction, according to the introduced model, one would expect a purely exponential decay of the expectation values with the number of process repetitions due to the Pauli noise. From this exponential decay, we can then characterize the Pauli error rates [7, 24] and mitigate these noise components using PEC [8]. If the final measurement statistics after coherent error mitigation and PEC closely follow the ideal statistics, we demonstrated that the proposed model is physical, and the characterization protocol provides valuable insights.
@@
Any unitary channel is reversible. Therefore, in our model, the effect of the coherent noise can be canceled by single- and two-qubit rotations with opposite rotation angles than the characterized ones. We introduce a parameterized circuit element (see Fig. 7) capable of representing all needed rotations while corresponding to the identity if all parameters are set to zero (see App. F). We use this circuit element to run echo-experiments for three different situations: unmitigated; coherent-error mitigated; and coherent-error mitigated combined with PEC. The results are presented in Fig. 4. Fig. 4(a) and (b) were obtained with 2048 shots per circuit, whereas for Fig. 4(c), an ensemble of 280 altered circuits with 512 shots each was executed per experiment.
@@
For the unmitigated case in Fig. 4(a), we see two main characteristics: Firstly, the expectation values of the 12 Pauli operators that should be ideally 00 are spreading. Secondly, the expectation values that ideally equal $1$, meaning that the states are eigenstates of the corresponding Pauli, decay. Note that the expectation value of the identity in the plot is not based on actual experiments and is always $1$ due to the trace constraints on density matrices.
@@
Applying the presented coherent noise characterization protocol on the data of Fig. 4(a) leads to a first estimate of the rotation angles of the coherent errors $\hat{\theta}_{0}$. To further improve the characterization, we run the characterization scheme with $\vec{\theta}=\hat{\theta}_{0}$ and obtain $\hat{\theta}_{1}$. This iterative approach is expected to improve the characterization because it compensates for violations of the assumed commutation of single- and two-qubit rotation errors. Fig. 4(b) shows the result of the final mitigation iteration with the estimate $\vec{\theta}=\hat{\theta}_{1}$ using the gate-based mitigation circuit from Fig. 7. The observed changes in the angles between the first and the last iteration were about an order of magnitude smaller than the biggest coherent errors and thus supporting the validity of the taken approximation. In Fig. 4(b), the mitigation effect is clearly visible, as the spreading behavior of the expectation values that are ideally 00 is largely suppressed. The expectation values of the Pauli operators, of which the prepared state is a $+1$ eigenstate, are higher in the mitigated case compared to the unmitigated estimation. However, the coherent noise mitigation does not retrieve their ideal-case expectation value of $1$, as according to our model, the statistic is still affected by Pauli noise.
@@
Figure 5: Influence of the coherent error mitigation on fidelity measurement between the ideal state and states evolved by the noisy identity. From the data of Fig. 8, the density matrix of the evolved state is estimated for each number of repetitions of the noisy identity and each of the 9 prepared states. The plot shows the fidelity between these state estimates and the ideally prepared state. The thin lines represent the evolution of the fidelity with the number of repetitions of the noisy identity for each state, and the thick lines are the average of the 9 thin lines.
@@
To show that the examined noise follows the coherently rotated Pauli noise model in Eq. (2), we estimate the Pauli error rates from the mitigated curves of Fig. 4(b) and apply PEC to mitigate the Pauli noise. Note that different from the above presented coherent noise correction, PEC is a mitigation technique that does not correct deterministically individual circuits (as Pauli channels are not reversible) but allows noise-free estimation of expectation values from circuit ensembles. It is important to emphasize that compared to previous works demonstrating PEC [58, 24], we do not use Pauli twirling and, thus, avoid the associated circuit and sampling overhead. The Pauli eigenvalues of the Pauli noise channel are estimated by fitting exponential curves to the expectation values that are supposed to be one in the ideal case. The Pauli error rates can then be found by a Walsh-Hadamard transform [7]. We follow the PEC protocol of van den Berg et al. [24] to mitigate the incoherent noise. It consists of applying the non-physical inverse of a Pauli channel by running an ensemble of unitary processes composing this inverse. Those results are weighted according to the Pauli error rates and combined in post-processing. The results of the echo-experiments with coherent error mitigation and PEC are displayed in Fig. 4(c). Compared to Fig. 4(b), the data are noisier, which we suspect to be due to the PEC sampling overhead.
@@
The obtained measurement statistics clearly resemble the main features of the ideal measurement statistics. Therefore, we can conclude that the introduced noise model explains the leading noise terms in the examined system and that the presented characterization protocol generates valid physical insights into the noise. The data of other input states is shown in App. G. Similar characterization and mitigation results were obtained on ibm_jakarta.
@@
Based on the data from Fig 4(a) and (b), we can reconstruct the density matrices for the $9$ prepared initial states at every repetition of the noisy identity. We then estimate their fidelity with their ideal initial state. The standard state fidelity formula is used $F(\sigma,\rho)=\mathrm{tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}$. We observe in Fig. 5 that the mitigation decreases the average infidelity by about a factor of two.
@@ V Generalization of the protocol
Furthermore, we can readily extend the proposed model to qudit-based quantum processors, which are also susceptible to coherent errors [59, 60, 61, 62, 63]. In these systems, Pauli $X$ and $Z$ operators are generalized as ${X_{d}\ket{s}=\ket{s+1\,\,\mathrm{mod}\,\,d}}$ and ${Z_{d}\ket{s}=\omega^{s}\ket{s}}$, respectively. Here, |$\ket{s}$ represents a qudit computational state with $s\in\{1,2,\ldots,d\}$ and $\omega=\exp(i2\pi/d)$. Consequently, the Pauli noise model for a $n$-qudit system can be expressed similarly to Eq. (1) but with $P_{i}\in\{(X_{d})^{p}(Z_{d})^{q}\,\,|\,\,q,p\in\{1,2,\ldots,d\}\}^{\otimes n}$ representing the generalized Pauli operators. Since the generalized Pauli operators exhibit similar properties to qubit Pauli gates (such as being unitary 1-designs and normalized by the qudit Clifford group), the Pauli transfer matrix formalism remains valid in the qudit framework. It thus enables the same twirling and characterization techniques employed in the qubit case to be utilized [59, 60, 61, 62, 63]. Moreover, this implies that the unitary errors can be effectively modeled by the coherently rotated Pauli noise model described in Eq. (2) and that the characterization protocol presented in Sec. III can be utilized. Additionally, the expansion of the protocol to generalized Pauli operators is also interesting for the characterization of coherent leakage of the qubit. Coherent leakage involves a unitary interaction between the computational subspace and the neglected part of the system’s Hilbert space or its environment [64]. Leakage back to the system introduces noise that is correlated to a past state of the system, introducing memory effects [65]. By expanding the system description to states in the leakage subspace, it is possible to describe the coherent leakage processes equally to coherent noise within the computational subspace. However, this demands precise control or generalized measurements of the leakage subspace, which must be clearly defined and localized.
@@ VI Conclusion and Outlook
In this paper, we introduced a characterization technique to estimate coherent errors in noisy quantum devices. An existing Pauli noise model is extended to incorporate coherent errors by adding unitary rotations to the model. We devise a procedure to learn the unknown parameters of this coherently rotated Pauli noise model. The scalability of our approach is affirmed by assuming a limited correlation length of the coherent noise. As the results of the characterization of a physical quantum processor showed, the protocol is feasible to be executed on an actual device. For the tested device, the characterization revealed the leading coherent errors introduced by the probed gate layer. Notably, these characterized coherent errors remained stable over multiple re-calibration cycles, suggesting a systematic nature. To verify the validity of our noise model, we mitigated the coherent errors and the characterized Pauli noise. The coherent error mitigation was achieved with a gate-level correction scheme for two-qubit coherent errors. To counter the Pauli noise, we applied probabilistic error cancellation. The combination of the two techniques resulted in almost ideal measurement statistics. This finding shows that the model is sufficiently expressible to explain the main noise contributions of a specific process on an actual processor. Moreover, the successful mitigation of the characterized errors confirms the applicability of the developed protocol. Our protocol is hardware agnostic and thus expected to be executable on a variety of hardware platforms.
Our work offers several promising avenues for future research. The method can enhance mitigation schemes that rely on detailed knowledge of the noise structure. One such approach is optimizing the resilience to coherent noise during circuit transpilation based on the noise insights for each gate or gate layer [68, 69, 70, 71, 72, 71]. A clear example of this is the hidden inverses method [69, 70], which mitigates the impact of coherent errors by compiling circuits in a way that induces destructive interference among the coherent errors of the gates. However, implementing hidden inverses requires knowledge of the physical nature of these errors, which in Ref. [70] was obtained through gate-set tomography. Integrating our approach into the calibration process can provide valuable information to the compiler, enabling it to leverage the hidden inverses technique while actively inverting the non-corrected coherent errors. Similarly, our technique opens the door to studying and developing noise-aware versions of advanced mitigation techniques such as dynamical decoupling or composite pulses. Moreover, information on coherent errors facilitates Pauli conjugation [73]. An efficient noise twirling method that limits sampling overhead compared to common Pauli twirling and, therefore, diminishes the reduction of the threshold of error-correcting codes.
@@
In the long run, the presented characterization protocol paves the way for the development of improved hardware platforms by identifying the limiting imperfections. Yet, already in the near term it can serve as a tool for various applications that rely on knowledge of the existing noise structure, such as quantum error correction or hardware-aware algorithm design.
@@ Appendix A Second order approximation in $\theta$ for PTM elements
To find the PTM element of $U_{\theta}$ up to the second order of $\theta$, we first expand $U_{\theta}=\Pi_{k}\text{exp}(-i\theta_{k}P_{k})$ using a second order approximation of Eq. (3):
$U_{\theta}$ $\approx\prod_{k}\left(1-\frac{\theta_{k}^{2}}{2}\right)\mathbb{I}-i\theta_{k}P_{k}$ (6)
@@
Using that $P_{a}P_{k}P_{b}P_{k}=(-1)^{\langle k,b\rangle}P_{a}P_{b}$, where $\langle k,b\rangle$ is 00 if $P_{b}$ and $P_{k}$ commute and $1$ otherwise, and the cyclic property of the trace leads to:
@@
where all terms of order in $\theta$ larger than two are dropped. Finally, we reach the approximation of the PTM element of $U_{\theta}$ up to the second order in $\theta$:
@@ Appendix B Closed form expression for estimation
The right hand side of Eq. (4) can be written equivalently as:
@@
where we use Einstein’s notation for the summation over $k$ and $C$ is a tensor with coefficients $k,i$ and $j$. Applying the channel to an arbitrary state $\left|\rho\right)$ leads to the equation ${\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)_{i}=\left|\rho\right)_{i}+\theta_{k}\,C_{kij}\left|\rho\right)_{j}}$, again applying Einstein’s notation to sum over $k$ and $j$. Defining the matrix $(B_{\left|\rho\right)})_{ik}=C_{kij}\left|\rho\right)_{j}$, the estimation problem boils down to a linear inverse problem of form ${B_{\left|\rho\right)}\,\theta=\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)-\left|\rho\right)}$ where the vectors $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ can be experimentally estimated and $B_{\left|\rho\right)}$ is a function of $\left|\rho\right)$. However, as $B_{\left|\rho\right)}$ is singular and cannot be inverted. To overcome this we must estimate $B_{\left|\rho\right)}$, $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ for multiple states $\rho$. Doing this for $N$ states, we formulate the overdetermined system $\tilde{B}\theta=y$, where $\tilde{B}$ of size $15N\times 15$ is the concatenation of all matrices $B_{\left|\rho_{i}\right)}$ and $y$ the concatenation of all vectors $\left|\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\right)-\left|\rho_{i}\right)$. Finally, we estimate $\theta$ using the least square method,
@@
Figure 6: Coherent noise drift. We repeated the characterization result of Fig. 3 at $3$ different times. Subplot (a) shows the baseline data of Fig. 3. Subplot (b) displays the observed changes in the coherent error characterization over a time span of 19 days and subplot (c) exhibits the differences over (6) hours. The experiments were run on the $7$-qubit IBM Quantum Processor ibm_lagos [32] on September 22, 2022, and October 11, 2022.

### v3

@@ III Characterization
We conduct the characterization of the same device and gate layer as in Fig. 3 three times, with time differences between the runs being 6 hours and 19 days. The results available in App. E show that the leading coherent errors in the system do not change over time. The most significant single-qubit coherent errors change by no more than $10\%$ and the two-qubit coherent rotation errors by maximally $0.018$ radians. This result suggests that the platform suffers from systematic coherent errors not influenced by the daily recalibration of the system.
@@ IV Mitigation
In principle, the characterization results in Fig. 3 may originate from overfitting the nonideal process to our model. Indeed, it is not known whether the coherently rotated Pauli noise model in Eq. (2) is expressive enough to resemble the main characteristics of the probed IBMq process. To verify that the model captures the dominant noise contributions of a process on the probed device, we examine whether the output of the presented protocol facilitates significant noise mitigation
@@
Returning to the unmitigated circuits in Fig. 1, our goal is to suppress the observed oscillatory behavior in the echo experiment by correcting the characterized coherent errors at the circuit level. In the case of a successful correction, according to our model, one would expect a purely exponential decay of the expectation values with the number of process repetitions due to the Pauli noise. From this exponential decay, we can then characterize the Pauli error rates [7, 24] and mitigate these noise components using PEC [8]. If the final measurement statistics after coherent error mitigation and PEC closely follow the ideal statistics, we have demonstrated that the proposed model is physical, and the characterization protocol provides valuable insights.
@@
Any unitary channel is reversible. Therefore, in our model, the effect of the coherent noise can be canceled by single- and two-qubit rotations with opposite rotation angles to those characterized. We introduce a parameterized circuit element (see Fig. 7) capable of representing all required rotations while corresponding to the identity if all parameters are set to zero (see App. G). We use this circuit element to run echo experiments for three different situations: unmitigated; coherent-error mitigated; and coherent-error mitigated combined with PEC. The results are presented in Fig. 4. Fig. 4(a) and (b) were obtained with 2048 shots per circuit, whereas for Fig. 4(c), an ensemble of 280 altered circuits with 512 shots each was executed per experiment.
@@
For the unmitigated case in Fig. 4(a), we see two main characteristics. Firstly, the expectation values of the 12 Pauli operators that should ideally be 00 are spreading. Secondly, the expectation values that ideally equal $1$, meaning that the states are eigenstates of the corresponding Pauli, decay.
@@
Applying our coherent noise characterization protocol to the data of Fig. 4(a) leads to a first estimate of the rotation angles of the coherent errors $\hat{\theta}_{0}$. To further improve the characterization, we run the characterization scheme with $\vec{\theta}=\hat{\theta}_{0}$ and obtain $\hat{\theta}_{1}$. This iterative approach is expected to improve the characterization because it compensates for violations of the assumed commutation of single- and two-qubit rotation errors. Fig. 4(b) shows the result of the final mitigation iteration with the estimate $\vec{\theta}=\hat{\theta}_{1}$ using the gate-based mitigation circuit from Fig. 7. The observed changes in the angles between the first and the last iteration were about an order of magnitude smaller than the biggest coherent errors and thus support the validity of the approximation taken. In Fig. 4(b), the mitigation effect is clearly visible, as the spreading behavior of the expectation values that are ideally 00 is largely suppressed. The expectation values of the Pauli operators, of which the prepared state is a $+1$ eigenstate, are higher in the mitigated case compared to the unmitigated estimation. However, the coherent noise mitigation does not retrieve their ideal-case expectation value of $1$, as according to our model the statistic is still affected by Pauli noise.
@@
Figure 5: Influence of coherent error mitigation on fidelity measurement between the ideal state and states evolved by the noisy identity. From the data of Fig. 8, the density matrix of the evolved state is estimated for each number of repetitions of the noisy identity and each of the nine prepared states. The plot shows the fidelity between these state estimates and the ideally prepared state. The thin lines represent the evolution of the fidelity with the number of repetitions of the noisy identity for each state, and the thick lines are the average of the nine thin lines.
@@
To show that the noise examined follows the coherently rotated Pauli noise model in Eq. (2), we estimate the Pauli error rates from the mitigated curves of Fig. 4(b) and apply PEC to mitigate the Pauli noise. Note that, different from the above-presented coherent noise correction, PEC is a mitigation technique that does not correct individual circuits deterministically (as Pauli channels are not reversible) but allows noise-free estimation of expectation values from circuit ensembles. It is important to emphasize, that compared to previous works demonstrating PEC [58, 24], we do not use Pauli twirling and thus avoid the associated circuit and sampling overhead. The Pauli eigenvalues of the Pauli noise channel are estimated by fitting exponential curves to the expectation values that are supposed to be 1 in the ideal case. The Pauli error rates can then be found by a Walsh-Hadamard transform [7]. We follow the PEC protocol of van den Berg et al. [24] to mitigate the incoherent noise. This involves applying the nonphysical inverse of a Pauli channel by running an ensemble of unitary processes composing this inverse. Those results are weighted according to the Pauli error rates and combined in postprocessing. The results of the echo experiments with coherent error mitigation and PEC are displayed in Fig. 4(c). Compared to Fig. 4(b), the data are noisier, which we suspect to be due to the PEC sampling overhead.
@@
The measurement statistics obtained clearly resemble the main features of the ideal measurement statistics. Therefore, we can conclude that our noise model explains the leading noise terms in the system examined and that the characterization protocol presented generates valid physical insights into the noise. The data of other input states is shown in App. H. Similar characterization and mitigation results were obtained on ibm_jakarta.
@@
Based on the data from Fig 4(a) and (b), we can reconstruct the density matrices for the nine prepared initial states at every repetition of the noisy identity. We then estimate their fidelity with their ideal initial state. The standard state fidelity formula is used $F(\sigma,\rho)=\mathrm{tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}$. We observe in Fig. 5 that the mitigation decreases the average infidelity by about a factor of 2.
@@ V Generalization of the protocol
Furthermore, we can readily extend the proposed model to qudit-based quantum processors, which are also susceptible to coherent errors [59, 60, 61, 62, 63]. In these systems, Pauli $X$ and $Z$ operators are generalized as ${X_{d}\ket{s}=\ket{s+1\,\,\mathrm{mod}\,\,d}}$ and ${Z_{d}\ket{s}=\omega^{s}\ket{s}}$, respectively. Here, |$\ket{s}$ represents a qudit computational state with $s\in\{1,2,\ldots,d\}$ and $\omega=\exp(i2\pi/d)$. Consequently, the Pauli noise model for an $n$-qudit system can be expressed similarly to Eq. (1) but with $P_{i}\in\{(X_{d})^{p}(Z_{d})^{q}\,\,|\,\,q,p\in\{1,2,\ldots,d\}\}^{\otimes n}$ representing the generalized Pauli operators. Since the generalized Pauli operators exhibit similar properties to qubit Pauli gates (such as being unitary 1-designs and normalized by the qudit Clifford group), the Pauli transfer matrix formalism remains valid in the qudit framework. It thus enables the same twirling and characterization techniques employed in the qubit case to be utilized [59, 60, 61, 62, 63]. Moreover, this implies that the unitary errors can be effectively modeled by the coherently rotated Pauli noise model described in Eq. (2) and that the characterization protocol presented in Sec. III can be utilized. Additionally, the expansion of the protocol to generalized Pauli operators is also of interest for the characterization of coherent leakage of the qubit. Coherent leakage involves a unitary interaction between the computational subspace and the neglected part of the system’s Hilbert space or its environment [64]. Leakage back to the system introduces noise that is correlated to a past state of the system, introducing memory effects [65]. By expanding the system description to states in the leakage subspace, it is possible to describe the coherent leakage processes equally to coherent noise within the computational subspace. However, this demands precise control or generalized measurements of the leakage subspace, which must be clearly defined and localized.
@@ VI Conclusion and Outlook
In this paper, we introduced a characterization technique to estimate coherent errors in noisy quantum devices. An existing Pauli noise model is extended to incorporate coherent errors by adding unitary rotations to the model. We devise a procedure to learn the unknown parameters of this coherently rotated Pauli noise model. The scalability of our approach is affirmed by assuming a limited correlation length of the coherent noise. As the results of the characterization of a physical quantum processor showed, it is feasible to execute the protocol on an actual device. For the device tested, the characterization revealed the leading coherent errors introduced by the probed gate layer. Notably, these characterized coherent errors remained stable over multiple recalibration cycles, suggesting a systematic nature. To verify the validity of our noise model, we mitigated the coherent errors and the characterized Pauli noise. Coherent error mitigation was achieved with a gate-level correction scheme for two-qubit coherent errors. To counter the Pauli noise, we applied probabilistic error cancelation. The combination of the two techniques resulted in almost ideal measurement statistics. This finding shows that the model is sufficiently expressible to explain the main noise contributions of a specific process on an actual processor. Moreover, the successful mitigation of the characterized errors confirms the applicability of the developed protocol. Our protocol is hardware agnostic and thus expected to be executable on a variety of hardware platforms.
Our work offers several promising avenues for future research. The method can enhance mitigation schemes that rely on detailed knowledge of the noise structure. One such approach is optimizing the resilience to coherent noise during circuit transpilation based on the noise insights for each gate or gate layer [68, 69, 70, 71, 72, 71]. A clear example of this is the hidden inverses method [69, 70], which mitigates the impact of coherent errors by compiling circuits in a way that induces destructive interference among the coherent errors of the gates. However, implementing hidden inverses requires knowledge of the physical nature of these errors, which in Ref. [70] was obtained through gate-set tomography. Integrating our approach into the calibration process can provide valuable information to the compiler, enabling it to leverage the hidden inverses technique while actively inverting the noncorrected coherent errors. Similarly, our technique opens the door to studying and developing noise-aware versions of advanced mitigation techniques such as dynamical decoupling or composite pulses. Moreover, information on coherent errors facilitates Pauli conjugation [73]. An efficient noise twirling method that limits sampling overhead compared to common Pauli twirling and, therefore, diminishes the reduction of the threshold of error-correcting codes.
@@
In the long run, the characterization protocol presented paves the way for the development of improved hardware platforms by identifying the limiting imperfections. Yet, already in the near term it can serve as a tool for various applications that rely on knowledge of the existing noise structure, such as quantum error correction or hardware-aware algorithm design.
@@ Acknowledgements.
Data Availability
The data supporting the findings of this study, including simulation and experimental results, are not publicly available but can be obtained from the corresponding author upon reasonable request.
@@ Appendix A Second order approximation in $\theta$ for PTM elements
To find the PTM element of $U_{\theta}$ up to the second-order of $\theta$, we first expand $U_{\theta}=\Pi_{k}\text{exp}(-i\theta_{k}P_{k})$ using a second order approximation of Eq. (3):
$U_{\theta}$ $\approx\prod_{k}\left(\left(1-\frac{\theta_{k}^{2}}{2}\right)\mathbb{I}-i\theta_{k}P_{k}\right)$ (6)
@@
Using that $P_{a}P_{k}P_{b}P_{k}=(-1)^{\langle k,b\rangle}P_{a}P_{b}$, where $\langle k,b\rangle$ is 00 if $P_{b}$ and $P_{k}$ commute and $1$ otherwise, and the cyclic property of the trace leads to
@@
where all terms of order in $\theta$ greater than 2 are dropped. Finally, we obtain the approximation of the PTM element of $U_{\theta}$ up to the second order in $\theta$:
@@ Appendix B Closed form expression for estimation
The right-hand side of Eq. (4) can be written equivalently as
@@
where we use Einstein’s notation for the summation over $k$ and $C$ is a tensor with coefficients $k,i$, and $j$. Applying the channel to an arbitrary state $\left|\rho\right)$ leads to the equation ${\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)_{i}=\left|\rho\right)_{i}+\theta_{k}\,C_{kij}\left|\rho\right)_{j}}$, again applying Einstein’s notation to sum over $k$ and $j$. Defining the matrix $(B_{\left|\rho\right)})_{ik}=C_{kij}\left|\rho\right)_{j}$, the estimation problem boils down to a linear inverse problem of the form ${B_{\left|\rho\right)}\,\theta=\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)-\left|\rho\right)}$ where the vectors $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ can be experimentally estimated and $B_{\left|\rho\right)}$ is a function of $\left|\rho\right)$. However, $B_{\left|\rho\right)}$ is singular and cannot be inverted. To overcome this we must estimate $B_{\left|\rho\right)}$, $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ for multiple states $\rho$. Doing this for $N$ states, we formulate the overdetermined system $\tilde{B}\theta=y$, where $\tilde{B}$ of size $15N\times 15$ is the concatenation of all matrices $B_{\left|\rho_{i}\right)}$ and $y$ the concatenation of all vectors $\left|\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\right)-\left|\rho_{i}\right)$. Finally, we estimate $\theta$ using the least-squares method,
@@
Figure 6: Coherent noise drift. We repeated the characterization result of Fig. 3 at three different times. Subplot (a) shows the baseline data of Fig. 3. Subplot (b) displays the observed changes in the coherent error characterization over a time span of 19 days and subplot (c) exhibits the differences over 6 hours. The circuits were executed on the seven-qubit IBM Quantum processor ibm_lagos [32] (specifications in App. F) on September 22, 2022 and October 11, 2022.